OpenStudy (anonymous):

factor this a2 – a – 20

OpenStudy (nory):

You always want to see first if there is a common factor. Is there?

OpenStudy (anonymous):

can you show your work too please

OpenStudy (ajprincess):

Can you pls tell me two factors of -20 b nd csuch that their sum is -1. i.e b+c=-1 b*c=-20

OpenStudy (anonymous):

-10 , -2

OpenStudy (ajprincess):

-10+(-2)=-10-2=-12

OpenStudy (ajprincess):

nd also -10*-2=20

OpenStudy (ajprincess):

so they are nt the factors u need

OpenStudy (ajprincess):

Any other factors? @allie_bear22

OpenStudy (nathan917):

(a-5) * (a+4) Would this be the answer?